Circular No. 8134

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2003ef, 2003eg, 2003eh, 2003ei, 2003ej, 2003ek Further to IAUC 8131, M. Ganeshalingam, J. Graham, H. Pugh, and

W. Li report the discovery of several supernovae on unfiltered KAIT images:

SN  2003 UT  α_{2000}  δ_{2000}  Mag.  Offset
2003eg May 17.3 12^{h}50^{m}58.^{s}33 −14^{o}20′01′′.1 15.8 16′′.7 E, 3′′.2 S
2003eh May 17.2 11 08 24.34 + 3 29 47.0 17.6 5′′.8 E, 5′′.6 S
2003ei May 19.4 16 28 40.30 +12 46 05.8 17.2 4′′.6 W, 9′′.5 N
2003ej May 20.3 12 39 11.10 + 0 43 30.4 16.5 4′′.0 W, 28′′.2 N
2003ek May 20.5 17 18 32.50 +41 40 00.8 18.2 3′′.6 E, 5′′.1 S

The offset of SN 2003ei is from the northern component of UGC 10402. Additional approximate magnitudes: SN 2003eg in NGC 4727, Apr. 14.3 UT, [18.5; May 18.3, 15.8. SN 2003eh in MCG +01-29-3, May 1.2, [18.5; 18.3, 17.6. SN 2003ei, Mar. 31.5, [19.5; May 20.4, 16.9. SN 2003ej in UGC 7820, May 11.3, [18.5; 21.3, 16.3. SN 2003ek, May 12.5, [19.0; 21.3, 18.2.

Further to IAUC 8129, T. Matheson, P. Challis, and R. Kirshner report on spectroscopy (range 370–750 nm) obtained by M. Calkins with the Mt. Hopkins 1.5-m telescope. A spectrum of SN 2003ef (cf. IAUC 8131), obtained on May 21.19 UT, shows it to be a type-II supernova, displaying a continuum with well-developed hydrogen P-Cyg lines; adopting the NASA/IPAC Extragalactic Database (NED) host-galaxy recession velocity (4103 km/s), the expansion velocity derived from the minimum of the Hβ line is 7900 km/s. A spectrum of SN 2003eg (May 21.21), showing it to be a type-II supernova, consists of a blue continuum with incipient P-Cyg lines of hydrogen. The NED host-galaxy recession velocity is 7495 km/s, but narrow lines from an H II region superposed on the spectrum of the supernova indicate a recession velocity of 4360 km/s; adopting the latter value, the expansion velocity derived from the minimum of the Hβ line is 10000 km/s. A spectrum of SN 2003eh (May 21.17) shows it to be a reddened type-Ia supernova near maximum; there is a strong (EW ∼ 0.3 nm) Na I D absorption line at the velocity of the host galaxy. Adopting the NED hostgalaxy recession velocity (7651 km/s), the supernova expansion velocity is 10300 km/s for Si II (rest 635.5 nm). The spectral-feature age (Riess et al. 1997, A.J. 114, 722) of SN 2003eh is 2 ± 2 days before maximum light. A spectrum of SN 2003ei (May 21.23), showing it to be a type-IIn supernova, displays a blue continuum with both broad and narrow hydrogen emission lines; there are weak lines that are consistent with Wolf-Rayet features, such as were seen in SN 1998S (e.g., Leonard et al. 2000, Ap.J. 536, 239).

2003 May 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green